# A Review of Large Language Models and Autonomous Agents in Chemistry

### v1

@@ 3.2.2 Encoder-only Mol-LLMs
In June 2023, Yuksel et al. 176 introduced SELFormer, building on ideas from ChemBERTa2177 and using SELFIES for large data input. Yuksel et al. 176 argue that SMILES strings have validity and robustness issues, hindering effective chemical interpretation of the data. SELFormer uses SELFIES and is pretrained on two million drug-like compounds, fine-tuned for diverse molecular property prediction tasks (BBBP, SIDER, Tox21, HIV, BACE, FreeSolv, ESOL, PDBbind from MoleculeNet).55 SELFormer outperformed all competing methods for some tasks and produced comparable results for the rest. It could also discriminate molecules with different structural properties. The paper suggests future directions in multimodal models combining structural data with other types of molecular information, including text-based annotations. We will discuss such multimodal models below.
@@
Also in 2023, Yu et al. 185 published SolvBERT, a multi-task BERT-based regression model that could predict both solvation free energy and solubility from the SMILES notations of solute-solvent complexes. It was trained on the CombiSolv-QM dataset,224 a curation of experimental solvent free energy data called CombiSolv-Exp-8780,225, 226, 227, 228 and the solubility dataset from Boobier et al. 186. SolvBERT’s performance was benchmarked against advanced graph-based models229, 230 This work is powerful because there is an expectation that solvation free energy depends on 3-dimensional conformational properties of the molecules, or at least 2D properties that would be well characterized by graph-based molecular representations. SolvBERT showed comparable performance to DMPNN in predicting solvation free energy, largely due to its effective clustering feature in the pretraining phase as shown by TMAP visualizations. Furthermore, SolvBERT outperformed GROVER in predicting experimentally evaluated solubility data for new solute-solvent combinations. This underscores the significance of SolvBERT’s ability to capture the dynamic and spatial complexities of solvation interactions in a text-based model.
@@
Despite significant progress, Jiang et al. 231 highlighted the limitations of labeled data in 2024 and introduced INTransformer,231 a transformer-based method for predicting molecule properties. INTransformer enhances the capture of global molecular information by adding perturbing noise and employing contrastive learning to artificially augment smaller datasets. This method achieved higher performance, and ongoing work continues to explore various transformer strategies for smaller datasets.
@@ 3.3 Property Directed Inverse Design and Decoder-only mol-LLMs
Table 2: Decoder-only scientific LLMs. aa:“Model Size” is reported as the number of parameters. “PubMed” refer to the PubMed abstracts dataset, while PMC (PubMed Corpus) refers to the full-text corpus dataset. $b$: The total number of parameters was not reported.
@@
Tx-LLM232 $b$ TDC datasets PaLM-2 Property prediction and retrosynthesis 2024.06
BioMedLM147 2.7B PubMed abstracts and full articles GPT QA 2024.03
BioMistral148 7B PubMed Central (PMC) Mistral QA 2024.02
BiMediX149 8x7B 1.3M Arabic-English instructions (BiMed) Mixtral QA 2024.02
EpilepsyLLM150 7B Data from the Japan Epilepsy Association, Epilepsy Information Center, and Tenkan Net LLaMa QA 2024.01
ChemSpaceAL233 $b$ ChEMBL 33, GuacaMol v1, MOSES, and BindingDB 08-2023 GPT Molecule Generation 2024.01
CheXagent151 7B 28 publicly available datasets, including PMC, MIMIC, wikipedia, PadChest, and BIMCV-COVID-19 Mistral QA, Image understanding 2024.01
BioMedGPT-LM152 7B and 10B 5.5M biomedical papers from S2ORC LLaMA2 QA 2023.08
Darwin153 7B SciQ and Web of Science LLaMA QA, Property prediction, NER, and Molecule Generation 2023.08
cMolGPT44 $b$ MOSES GPT Molecule Generation 2023.04
PMC-LLaMA154 7B and 13B MedC-k and MedC-I LLaMA QA 2023.04
Regression Transformer173 $\sim$27M ChEMBL, MoleculeNet, USPTO, etc XLNet Property prediction, Molecule tunning, Molecule generation 2023.04
Galactica155 125M, 1.3B, 6.7B, 30B, 120B The galactica corpus, a curation with 62B scientific documents Decoder-only QA, NER, Document Summarization, Property Prediction 2022.11
BioGPT156 355M 15M of Title and abstract from PubMed GPT-2 QA, NER, and Document Classification 2022-09
MolGPT234 6M MOSES and GuacaMol GPT Molecule Generation 2021.10
Adilov2021235 13.4M 5M SMILES from ChemBERTa’s PubChem-10M. GPT-2 Property prediction and molecule generation 2021.09
Decoder-only GPT-like architectures offer significant value for property directed molecule generation and de novo chemistry applications because they excel in generating novel molecular structures by learning from vast datasets of chemical compounds, and capture intricate patterns and relationships within molecular sequences. They thus propose viable new compounds that adhere to desired chemical properties and constraints, enabling rapid exploration and innovation within an almost infinite chemical space. We first describe property directed inverse design from a chemistry perspective, and then go on to provide examples of how decoder-only LLMs have propelled inverse design forward.
@@ 3.3.1 Property Directed Inverse Design
Nature has been a significant source of molecules that inhibit disease proliferation, as organisms have evolved chemicals to protect themselves. This has led to most pharmaceuticals being derived from natural products,236, 237 which offer advantages such as cell permeability, target specificity, and a vast chemical diversity.238 However, despite these benefits, there are high costs and complexities associated with high-throughput screening and synthesis of natural products.238, 236 While building on the chemical diversity and efficacy of natural products, AI also opens new avenues for synthesizing unique compounds efficiently. Advanced in-silico molecular design allows for rapid mutation239 towards valid de-novo molecular structures that are synthesizable,211, 240 streamlining the iterative process of drug development. Yet, the true innovation lies in the potential of LLM-driven target-directed molecular design or “Inverse Design,” where we can start with a desired property and directly generate molecules that manifest this attribute, bypassing traditional stepwise modifications.241 This capability dramatically accelerates the pathway from concept to viable therapeutic agents and aligns well with decoder-only LLM architectures.
@@ 3.3.2 Decoder-only mol-LLMs
In 2021, Adilov 235 presented “Generative pretraining from Molecules,” one of the first applications of decoder-only models to SMILES strings. It pretrained a GPT-2-like causal transformer for self-supervised learning, introducing “adapters” between attention blocks for task-specific fine-tuning.242 This method, requiring minimal architectural changes, offered versatility in molecule generation and property prediction, aiming to surpass ChemBERTa’s encoder-only performance with a more scalable and resource-efficient approach. Another early decoder-only SMILES-based model was Bagal et al’s MolGPT model,234. MolGPT, with a mere 6 million parameters, advanced GPT-type LLMs for molecular generation. Its decoder-only architecture with masked self-attention facilitated learning long-range dependencies, enabling chemically valid SMILES representations that met complex structural rules involving valency and ring closures. The paper also used salience measures for interpretability in predicting SMILES tokens. MolGPT outperformed many existing Variational Auto-Encoder (VAE) based approaches243, 244, 245, 246, 247, 248, 249, 250 in predicting novel molecules with specified properties, being trained on datasets like MOSES251 and GuacaMol.252 It showed good performance with metrics like validity, uniqueness, Frechet ChemNet Distance (FCD),253 and KL divergence,252. While MolGPT’s computational demands might be higher than traditional VAEs, its ability to generate high-quality, novel molecules justifies this trade-off, but future research could likely optimize model efficiency or explore lighter versions. A brief summary of advancements in transformer-based models for de-novo molecule generation from 2023 and 2024 follows.
@@
Haroon et al. 254 further developed a GPT-based model with relative attention for de novo drug design, showing improved validity, uniqueness, and novelty. Both Wang et al. 255 and Mao et al. 256 presented work that surpassed MolGPT, with Mao et al.’s T5-type model257 generating novel compounds using IUPAC names directly.258 Although T5-based, we include this here because of its relevance to de novo drug design. In a similar vein, Zhang et al. 259 proposed including target 3D structural information in molecular generative models, even though their approach is not LLM-based. However, we nonetheless note its value for future structure-based LLM drug design. They demonstrated that integrating additional biological data significantly enhances the relevance and specificity of generated molecules for targeted drug discovery. Wang et al. 260 discussed PETrans, a deep learning method for generating target-specific ligands using protein-specific encoding and transfer learning. This study emphasized the application of transformer models for generating molecules with a high binding affinity to specific protein targets.
@@
In 2024, Yoshikai et al. 261 discussed the limitations of transformer architectures in recognizing chirality from SMILES representations, highlighting the challenges in learning overall molecular structures with, in particular, chirality impacting the prediction accuracy of molecular properties. They coupled a transformer with a VAE to address this. Qian et al. 262 introduced CONSMI, motivated by contrastive learning in NLP, to generate new molecules using multiple SMILES representations, improving molecular novelty and validity. Kyro et al. 233 presented ChemSpaceAL, an active learning method for protein-specific molecular generation, efficiently discovering molecules with desired characteristics without prior knowledge of existing inhibitors. Yan et al. 263 proposed the GMIA framework, featuring a graph mutual interaction attention decoder for drug-drug interaction prediction, enhancing prediction accuracy and interpretability. Lastly, Shen et al. 264 reported on AutoMolDesigner, an AI-based open-source software for automated design of small-molecule antibiotics.
@@
For a deeper dive into decoder-only transformer architecture in chemistry, we highlight the May 2023 “Taiga” model by Mazuz et al. 85, and cMolGPT by Wang et al. 44. Taiga first learns to map SMILES strings to a vector space, and then refines that space using a smaller dataset of labeled molecules to produce molecules with targeted attributes. It employs an autoregressive mechanism, sequentially predicting each SMILES character based on the previous ones. For optimizing molecular properties, Taiga uses the REINFORCE algorithm90, which develops a strategy to enhance desired molecular features. Although this fine-tuning with reinforcement learning (RL) slightly reduces molecule validity, it significantly boosts their practical applicability. Initially evaluated using the Quantitative Estimate of Drug-likeness (QED) metric,265 Taiga has also demonstrated promising results in targeting IC50 values,133 the BACE protein,266 and anti-cancer activities they collected from a variety of sources. This work underscores the importance of using new models to address applications that require a higher level of chemical sophistication, to illustrate how such models could ultimately be applied outside of the available benchmark datasets. It also builds on the necessary use of standardized datasets and train-validation-test splitting, to demonstrate progress, as explained by Wu et al. 55. Yet, even the MoleculeNet benchmarks55 are flawed and we point the reader here to a more detailed discussion on benchmarking,142 given that a significant portion of molecules in the BACE dataset have undefined stereo centers, which, at a deeper level, complicates the modeling and prediction accuracy.
@@
Wang et al. 44 introduced “cMolGPT: A Conditional Generative Pre-Trained Transformer for Target-Specific de novo Molecular generation,” which underscores the importance of incorporating chemical domain knowledge to effectively navigate the vast landscape of drug-like molecules. Using unsupervised learning and an auto-regressive approach, cMolGPT generates SMILES guided by predefined conditions based on target proteins and binding molecules. Initially trained on the MOSES dataset251 without target information, the model is fine-tuned with embeddings of protein-binder pairs, focusing on generating compound libraries and target-specific molecules for the EGFR, HTR1A, and S1PR1 protein datasets.267, 268, 269, 270
@@
Their approach employs a QSAR model5 to predict the activity of generated compounds, achieving a Pearson correlation coefficient over 0.75. However, reliance on this QSAR model, which has its own limitations, underscores the necessity for more extensive experimental datasets. Despite its capabilities, the cMolGPT model tends to generate molecules within the confines of the same sub-chemical space as the original dataset, indicating that while it successfully identifies potential binders, it faces challenges in broadly exploring the chemical space to discover novel solutions. This limitation highlights the difficulty of generating molecules from significantly different chemical families that are effective in binding or possess a combination of desired properties, suggesting a need for approaches that can more effectively expand the exploration of chemical diversity. Regardless, models capable of this kind of conditional generation are typically trained on large datasets of known protein-ligand interactions. Through this training, they learn the complex relationship between protein structures and the chemical structures of binding ligands. The model learns to predict which molecular features or structures are important for interacting with proteins that have certain characteristics (as captured in the embeddings). While both “Taiga” and “cMolGPT” use transformer architectures and decoder-only setups for molecule generation, they differ in approach. Taiga integrates reinforcement learning to optimize its generative model, while cMolGPT uses target-specific embeddings to guide the generation process.
@@
Outside of training on SMILES strings, Mao et al. 258 introduce iupacGPT, which also uses a decoder-only architecture, modeled after GPT-2. It addresses the limitations of SMILES strings by using IUPAC names to integrate human-readable chemical semantics. It focuses on pretraining with a vast IUPAC dataset and fine-tuning with lightweight networks, excelling in molecule generation, classification, and regression tasks.
@@
We also refer the reader to several other good reviews, 271, 272, 273, 129 with Goel et al. 274 exploring the efficiency of modern machine learning methods in sampling drug-like chemical space for virtual screening and molecular design. The work of Goel et al. 274 discussed generative models, including LLM approaches, for approximating the entire drug-like chemical space and highlighted models conditioned on specific properties or receptor structures.
@@
We provide a segue from this section by introducing the work by Jablonka et al. 275, which showcases a decoder-only GPT model that, despite its training on natural language rather than specialized chemical languages, competes effectively with decoder-only LLMs tailored to chemical languages. The authors use GPT-3 to answer complex chemistry questions and therefore highlight its potential as a foundational tool in the field. This sets the stage for integrating a natural language decoder-only LLM, like ChatGPT, into chemical research, where it could serve as a central hub. This foreshadows future developments to pair the LLM with specialized tools to enhance capabilities, paving the way for the creation of autonomous agents that leverage deep language understanding in scientific domains.

### v2

@@ 3.2.2 Encoder-only Mol-LLMs
In June 2023, Yuksel et al. 149 introduced SELFormer, building on ideas from ChemBERTa2150 and using SELFIES for large data input. Yuksel et al. 149 argue that SMILES strings have validity and robustness issues, hindering effective chemical interpretation of the data, although this perspective is not universally held.235 SELFormer uses SELFIES and is pretrained on two million drug-like compounds, fine-tuned for diverse molecular property prediction tasks (BBBP, SIDER, Tox21, HIV, BACE, FreeSolv, ESOL, PDBbind from MoleculeNet).57 SELFormer outperformed all competing methods for some tasks and produced comparable results for the rest. It could also discriminate molecules with different structural properties. The paper suggests future directions in multimodal models combining structural data with other types of molecular information, including text-based annotations. We will discuss such multimodal models below.
@@
Also in 2023, Yu et al. 194 published SolvBERT, a multi-task BERT-based regression model that could predict both solvation free energy and solubility from the SMILES notations of solute-solvent complexes. It was trained on the CombiSolv-QM dataset,236 a curation of experimental solvent free energy data called CombiSolv-Exp-8780,237, 238, 239, 240 and the solubility dataset from Boobier et al. 195. SolvBERT’s performance was benchmarked against advanced graph-based models241, 242 This work is powerful because there is an expectation that solvation free energy depends on 3-dimensional conformational properties of the molecules, or at least 2D properties that would be well characterized by graph-based molecular representations. It shows an overachieving utility of using SMILES strings in property prediction, and aligns with other work by Winter et al. 243, regarding activity coefficients. SolvBERT showed comparable performance to DMPNN in predicting solvation free energy, largely due to its effective clustering feature in the pretraining phase as shown by TMAP visualizations. T Furthermore, SolvBERT outperformed GROVER in predicting experimentally evaluated solubility data for new solute-solvent combinations. This underscores the significance of SolvBERT’s ability to capture the dynamic and spatial complexities of solvation interactions in a text-based model.
@@
Despite significant progress, Jiang et al. 244 highlighted the limitations of labeled data in 2024 and introduced INTransformer,244 a transformer-based method for predicting molecule properties. INTransformer enhances the capture of global molecular information by adding perturbing noise and employing contrastive learning to artificially augment smaller datasets. This method achieved higher performance, and ongoing work continues to explore various transformer strategies for smaller datasets.
@@ 3.3 Property Directed Inverse Design and Decoder-only mol-LLMs
Table 2:  Decoder-only scientific LLMs. $a$:“Model Size” is reported as the number of parameters. “PubMed” refer to the PubMed abstracts dataset, while PMC (PubMed Corpus) refers to the full-text corpus dataset. $b$: The total number of parameters was not reported.
@@
Tx-LLM245 $b$ TDC datasets PaLM-2 Property prediction and retrosynthesis 2024.06
LlasMol246 $\sim$ 7B SMolInstruct Galactica, LLaMa, Mistral Property prediction, molecule captioning, molecule generation, retrosynthesis, name conversion 2024.04
BioMedLM118 2.7B PubMed abstracts and full articles GPT QA 2024.03
BioMistral119 7B PubMed Central (PMC) Mistral QA 2024.02
BiMediX120 8x7B 1.3M Arabic-English instructions (BiMed) Mixtral QA 2024.02
EpilepsyLLM121 7B Data from the Japan Epilepsy Association, Epilepsy Information Center, and Tenkan Net LLaMa QA 2024.01
ChemSpaceAL247 $b$ ChEMBL 33, GuacaMol v1, MOSES, and BindingDB 08-2023 GPT Molecule Generation 2024.01
CheXagent122 7B 28 publicly available datasets, including PMC, MIMIC, wikipedia, PadChest, and BIMCV-COVID-19 Mistral QA, Image understanding 2024.01
ChemGPT248 $\sim$ 1B 10M molecules from PubChem GPT-neo Molecule generation 2023.11
BioMedGPT-LM123 7B and 10B 5.5M biomedical papers from S2ORC LLaMA2 QA 2023.08
Darwin124 7B SciQ and Web of Science LLaMA QA, Property prediction, NER, and Molecule Generation 2023.08
cMolGPT46 $b$ MOSES GPT Molecule Generation 2023.04
PMC-LLaMA125 7B and 13B MedC-k and MedC-I LLaMA QA 2023.04
Regression Transformer145 $\sim$27M ChEMBL, MoleculeNet, USPTO, etc XLNet Property prediction, Molecule tuning, Molecule generation 2023.04
GPTChem249 175B Curation of multiple classification and regression benchmarks GPT-3 Property prediction and inverse design 2023.02
Galactica126 125M, 1.3B, 6.7B, 30B, 120B The galactica corpus, a curation with  62B scientific documents Decoder-only QA, NER, Document Summarization, Property Prediction 2022.11
BioGPT127 355M 15M of Title and abstract from PubMed GPT-2 QA, NER, and Document Classification 2022-09
SPT243 6.5M Synthetic data generated with the thermodynamic model COSMO-RS GPT-3 Property prediction 2022.09
MolGPT250 6M MOSES and GuacaMol GPT Molecule Generation 2021.10
Adilov2021251 13.4M 5M SMILES from ChemBERTa’s PubChem-10M. GPT-2 Property prediction and molecule generation 2021.09
Decoder-only GPT-like architectures offer significant value for property directed molecule generation and de novo chemistry applications because they excel in generating novel molecular structures by learning from vast datasets of chemical compounds, and capture intricate patterns and relationships within molecular sequences. They thus propose viable new compounds that adhere to desired chemical properties and constraints, enabling rapid exploration and innovation within an almost infinite chemical space. In addition, such large general-purpose models have been proven to enable specific applications by fine-tuning with a small amount of scientific data.249, 252 We first describe property-directed inverse design from a chemistry perspective and then provide examples of how decoder-only LLMs have propelled inverse design forward.
@@ 3.3.1 Property Directed Inverse Design
Nature has been a significant source of molecules that inhibit disease proliferation, as organisms have evolved chemicals to protect themselves. This has led to most pharmaceuticals being derived from natural products,253, 254 which offer advantages such as cell permeability, target specificity, and a vast chemical diversity.255 However, despite these benefits, there are high costs and complexities associated with high-throughput screening and synthesis of natural products.255, 253 While building on the chemical diversity and efficacy of natural products, AI also opens new avenues for synthesizing unique compounds efficiently. Advanced in-silico molecular design allows for rapid mutation256 towards valid de-novo molecular structures that are synthesizable,222, 257 streamlining the iterative process of drug development. Yet, the true innovation lies in the potential of LLM-driven target-directed molecular design or “Inverse Design,” where we can start with a desired property and directly generate molecules that manifest this attribute, bypassing traditional stepwise modifications.258 This capability dramatically accelerates the pathway from concept to viable therapeutic agents and aligns well with decoder-only LLM architectures.
@@ 3.3.2 Decoder-only mol-LLMs
In 2021, Adilov 251 presented “Generative pretraining from Molecules,” one of the first applications of decoder-only models to SMILES strings. It pretrained a GPT-2-like causal transformer for self-supervised learning, introducing “adapters” between attention blocks for task-specific fine-tuning.259 This method, requiring minimal architectural changes, offered versatility in molecule generation and property prediction, aiming to surpass ChemBERTa’s encoder-only performance with a more scalable and resource-efficient approach. Another early decoder-only SMILES-based model was Bagal et al’s MolGPT model,250. MolGPT, with a mere 6 million parameters, advanced GPT-type LLMs for molecular generation. Its decoder-only architecture with masked self-attention facilitated learning long-range dependencies, enabling chemically valid SMILES representations that met complex structural rules involving valency and ring closures. The paper also used salience measures for interpretability in predicting SMILES tokens. MolGPT outperformed many existing Variational Auto-Encoder (VAE) based approaches260, 261, 262, 263, 264, 265, 266, 267 in predicting novel molecules with specified properties, being trained on datasets like MOSES268 and GuacaMol.269 It showed good performance with metrics like validity, uniqueness, Frechet ChemNet Distance (FCD),270 and KL divergence,269. While MolGPT’s computational demands might be higher than traditional VAEs, its ability to generate high-quality, novel molecules justifies this trade-off, but future research could likely optimize model efficiency or explore lighter versions. A brief summary of advancements in transformer-based models for de-novo molecule generation from 2023 and 2024 follows.
@@
Haroon et al. 271 further developed a GPT-based model with relative attention for de novo drug design, showing improved validity, uniqueness, and novelty. Frey et al. 248 introduced ChemGPT and explored the challenges of hyperparameter tuning at scale in new domains, along with a consideration of how the scale of the pretraining dataset will influence neural architecture selection, relative to ”typical" chemical generative models. Both Wang et al. 272 and Mao et al. 273 presented work that surpassed MolGPT, with Mao et al.’s T5-type model274 generating novel compounds using IUPAC names directly.275 Although T5-based, we include this here because of its relevance to de novo drug design. In a similar vein, Zhang et al. 276 proposed including target 3D structural information in molecular generative models, even though their approach is not LLM-based. However, we nonetheless note its value for future structure-based LLM drug design. They demonstrated that integrating additional biological data significantly enhances the relevance and specificity of generated molecules for targeted drug discovery. Wang et al. 277 discussed PETrans, a deep learning method for generating target-specific ligands using protein-specific encoding and transfer learning. This study emphasized the application of transformer models for generating molecules with a high binding affinity to specific protein targets.
@@
In 2024, Yoshikai et al. 278 discussed the limitations of transformer architectures in recognizing chirality from SMILES representations, highlighting the challenges in learning overall molecular structures with, in particular, chirality impacting the prediction accuracy of molecular properties. They coupled a transformer with a VAE to address this. Qian et al. 279 introduced CONSMI, motivated by contrastive learning in NLP, to generate new molecules using multiple SMILES representations, improving molecular novelty and validity. Kyro et al. 247 presented ChemSpaceAL, an active learning method for protein-specific molecular generation, efficiently discovering molecules with desired characteristics without prior knowledge of existing inhibitors. Yan et al. 280 proposed the GMIA framework, featuring a graph mutual interaction attention decoder for drug-drug interaction prediction, enhancing prediction accuracy and interpretability. Lastly, Shen et al. 281 reported on AutoMolDesigner, an AI-based open-source software for automated design of small-molecule antibiotics.
@@
For a deeper dive into decoder-only transformer architecture in chemistry, we highlight the May 2023 “Taiga” model by Mazuz et al. 92, and cMolGPT by Wang et al. 46. Taiga first learns to map SMILES strings to a vector space, and then refines that space using a smaller dataset of labeled molecules to produce molecules with targeted attributes. It employs an autoregressive mechanism, sequentially predicting each SMILES character based on the previous ones. For optimizing molecular properties, Taiga uses the REINFORCE algorithm97, which develops a strategy to enhance desired molecular features. Although this fine-tuning with reinforcement learning (RL) slightly reduces molecule validity, it significantly boosts their practical applicability. Initially evaluated using the Quantitative Estimate of Drug-likeness (QED) metric,282 Taiga has also demonstrated promising results in targeting IC50 values,179 the BACE protein,283 and anti-cancer activities they collected from a variety of sources. This work underscores the importance of using new models to address applications that require a higher level of chemical sophistication, to illustrate how such models could ultimately be applied outside of the available benchmark datasets. It also builds on the necessary use of standardized datasets and train-validation-test splitting, to demonstrate progress, as explained by Wu et al. 57. Yet, even the MoleculeNet benchmarks57 are flawed, and we point the reader here to a more detailed discussion on benchmarking,186 given that a significant portion of molecules in the BACE dataset have undefined stereo centers, which, at a deeper level, complicates the modeling and prediction accuracy.
@@
Wang et al. 46 introduced “cMolGPT: A Conditional Generative Pre-Trained Transformer for Target-Specific de novo Molecular generation,” which underscores the importance of incorporating chemical domain knowledge to effectively navigate the vast landscape of drug-like molecules. Using unsupervised learning and an auto-regressive approach, cMolGPT generates SMILES guided by predefined conditions based on target proteins and binding molecules. Initially trained on the MOSES dataset268 without target information, the model is fine-tuned with embeddings of protein-binder pairs, focusing on generating compound libraries and target-specific molecules for the EGFR, HTR1A, and S1PR1 protein datasets.284, 285, 286, 287
@@
Their approach employs a QSAR model5 to predict the activity of generated compounds, achieving a Pearson correlation coefficient over 0.75. However, reliance on this QSAR model, which has its own limitations, underscores the necessity for more extensive experimental datasets. Despite its capabilities, the cMolGPT model tends to generate molecules within the confines of the same sub-chemical space as the original dataset, indicating that while it successfully identifies potential binders, it faces challenges in broadly exploring the chemical space to discover novel solutions. This limitation highlights the difficulty of generating molecules from significantly different chemical families that are effective in binding or possess a combination of desired properties, suggesting a need for approaches that can more effectively expand the exploration of chemical diversity. Regardless, models capable of this kind of conditional generation are typically trained on large datasets of known protein-ligand interactions. Through this training, they learn the complex relationship between protein structures and the chemical structures of binding ligands. The model learns to predict which molecular features or structures are important for interacting with proteins that have certain characteristics (as captured in the embeddings). While both “Taiga” and “cMolGPT” use transformer architectures and decoder-only setups for molecule generation, they differ in approach. Taiga integrates reinforcement learning to optimize its generative model, while cMolGPT uses target-specific embeddings to guide the generation process. Yu et al. 246 follow a different approach. LlaSMol246 starts from pretrained models (for instance Galactica, LlaMa2, and Mistral) and performs parameter efficient fine-tuning (PEFT)288, 289. Specifically, they used LoRa290. They show that PEFT has the potential to achieve state-of-the-art performance in property prediction when fine-tuning on MoleculeNet57.
@@
Outside of training on SMILES strings, Mao et al. 275 introduce iupacGPT, which also uses a decoder-only architecture, modeled after GPT-2. It addresses the limitations of SMILES strings by using IUPAC names to integrate human-readable chemical semantics. It focuses on pretraining with a vast IUPAC dataset and fine-tuning with lightweight networks, excelling in molecule generation, classification, and regression tasks.
@@
We also refer the reader to several other good reviews, 291, 292, 293, 175 with Goel et al. 294 exploring the efficiency of modern machine learning methods in sampling drug-like chemical space for virtual screening and molecular design. The work of Goel et al. 294 discussed generative models, including LLM approaches, for approximating the entire drug-like chemical space and highlighted models conditioned on specific properties or receptor structures.
@@
We provide a segue from this section by introducing the work by Jablonka et al. 249, which showcases a decoder-only GPT model that, despite its training on natural language rather than specialized chemical languages, competes effectively with decoder-only LLMs tailored to chemical languages. The authors use finetuned GPT-3 to answer complex chemistry questions and therefore highlight its potential as a foundational tool in the field. This sets the stage for integrating a natural language decoder-only LLM, like ChatGPT, into chemical research, where it could serve as a central hub. This foreshadows future developments to pair the LLM with specialized tools to enhance capabilities, paving the way for the creation of autonomous agents that leverage deep language understanding in scientific domains.